# Knowledge Transfer from Vision Foundation Models for Efficient Training of Small Task-specific Models

### v2

@@ 1 Introduction
Figure 1: Downstream task performance (EuroSAT classification with 10 labeled training images per class) of FastViT target model with different pretraining approaches. Task-oriented knowledge transfer from DINOv2 VFM with generic web data (CC3M) outperforms the popular ImageNet, CLIP and DINO pretraining approaches. Transfer set curated with image retrieval performs significantly better than generic web data.
Currently, the computer vision community is witnessing the emergence of various vision and multi-modal foundation models pretrained on massive datasets (Radford et al., 2021; Yuan et al., 2021; Kirillov et al., 2023; Oquab et al., 2023; Li et al., 2023b; Wang et al., 2023b). These models have been shown to work well for many downstream tasks, especially, when task-specific labeled data is limited (Radford et al., 2021). However, they cannot be used for many resource-constrained applications due to their high computational cost. Also, many applications such as autonomous driving, medical image diagnostics, and industrial automation, focus on specific tasks or domains and need small task-specific models rather than a large Vision Foundation Model (VFM). This raises an important, timely and yet underexplored question: How can we leverage the knowledge from a large VFM to effectively train a small task-specific model for a new target task with limited labeled training data?
@@ 2 Approach and contributions
We compare the proposed task-oriented knowledge transfer approach with task-agnostic knowledge transfer from VFMs, web-scale CLIP pretraining, supervised ImageNet pretraining, and self-supervised DINO (Caron et al., 2021) pretraining on five target tasks under limited labeled data settings using two VFMs, namely DINOv2-ViT-L/14 (Oquab et al., 2023) and OpenCLIP-ViT-L/14 (Ilharco et al., 2021), and two target mobile architectures, namely FastViT-S12 (Vasu et al., 2023a) and MobileViT-V2-1.0. We experiment with two transfer sets, a generic CC3M (Sharma et al., 2018) transfer set and a target task-related transfer set, and present several insightful findings to the community. To the best of our knowledge, there is no existing work that concretely establishes the below findings about leveraging VFMs for training small task-specific models with limited labeled data.
@@
  • Task-oriented knowledge transfer from VFMs outperforms the compute-intensive CLIP, widely-used supervised ImageNet and self-supervised DINO pretraining approaches by large margins (see Table 1), and is also computationally efficient (see Figures 1 and 3).
@@ 3.1 Experimental setup
Alternative approaches: IN-Pretrain: Supervised pretraining on 1.28M labeled images from the ImageNet-1K dataset (Deng et al., 2009). CLIP-Pretrain: Contrastive language image pretraining on 0.7B image-text pairs from (Cao et al., 2023) following the affinity mimicking CLIP distillation approach of (Wu et al., 2023)111 (Wu et al., 2023) has shown that affinity mimicking distillation works better than the CLIP loss of (Radford et al., 2021). We use the best ViT-B/16 model from (Mehta et al., 2022) as teacher. DINO-Pretrain: Self-supervised pretraining on the unlabeled transfer set following the DINO approach of (Caron et al., 2021).
@@
Target models: We use two recent efficient architectures, namely FastViT-S12 (Vasu et al., 2023a) and MobileViT-V2-1.0 (Mehta & Rastegari, 2023), as image encoders for the target models. For brevity, we refer to them as FastViT and MViT-V2, respectively. We present the results for FastViT in this section and the results for MViT-V2 in Section C.2.
Task-specific heads: We use a linear classifier for classification tasks, and a DeepLabV3 head (Chen et al., 2017) for segmentation tasks. See Section A.3 for further details.
We run each finetuning experiment three times and report the average results.
@@ 3.2.1 Performance improvement
When task-related transfer set is used, task-oriented transfer even outperforms finetuned VFM for some of the tasks when the amount of labeled data is small. This could be because the knowledge transfer process uses large-scale unlabeled target domain data that is not used while finetuning VFMs.
@@ 3.4 Retrieval augmented knowledge transfer
Specifically, we use the limited target task labeled dataset as the query set $\mathcal{Q}$ and a large set of images sourced from web as the gallery $\mathcal{G}$. We use an encoder network $\phi$ to map all the images to a $d$-dimensional embedding space, and perform retrieval based on Euclidean distances in this space following a query balanced approach. For a query image $x_{q}\in\mathcal{Q}$, we define $k$-NN$(x_{q})$ to be the set of its $k$ nearest neighbors from the gallery $\mathcal{G}$. To retrieve $N$ images in total, we find the smallest $k$ for which $\bigcup_{x_{q}\in\mathcal{Q}}k$-NN$(x_{q})$ contains at least $N$ unique images. If $\bigcup_{x_{q}\in\mathcal{Q}}k$-NN$(x_{q})$ contains more than $N$ images, we drop the $k^{th}$ neighbor of randomly selected queries until the retrieved set has $N$ images. By giving equal weight to all the queries, this approach increases the diversity of the retrieved samples. We combine the retrieved set with the initial query set and use this retrieval augmented transfer set for task-oriented knowledge transfer.
@@
3.4.1 ADE20K segmentation
Figure 5 (left) shows ADE20K segmentation performance of task-oriented knowledge transfer with retrieval augmented transfer sets of different sizes. Here, we use 4800 labeled images as both finetuning dataset and retrieval query set. The segmentation performance increases with the transfer set size until we reach 154K images and drops after that indicating that the size of YFCC15M subset that is most useful for ADE20K segmentation is around 154K. Using 154K as the target transfer set size, we curate different transfer sets by varying the number of query images used for retrieval. Table 3 (top) compares these curated transfer sets with the generic CC3M and the full ADE20K transfer sets. Retrieval augmented transfer sets clearly outperform both ADE20K and generic CC3M transfer sets.  [^3] The best performance in Figure 5(left) is lower than the performance for curated transfer set corresponding to 4800 query images in Table 3(top). This is because, we used shorter training runs (60K steps) to get the results in Figure 5(left), and once we identified the best transfer set size, we used longer training runs (200K steps) to get the results in Table 3(top).
@@ 4 Related works
Knowledge distillation is a widely-used approach for transferring knowledge between model architectures by training one model to mimic the outputs of another model. Numerous distillation approaches have been proposed over the past decade based on various knowledge representations such as task logits (Hinton et al., 2015), intermediate features or embeddings (Heo et al., 2019; Tian et al., 2020), relations between samples (Park et al., 2019), attention maps (Zagoruyko & Komodakis, 2017), etc. Please refer to Wang & Yoon (2022); Hu et al. (2023) for an overview of existing distillation approaches. Some recent works have specifically focused on multi-modal distillation of image-language models (Fang et al., 2021; Li et al., 2023c; Wang et al., 2023a; Sun et al., 2023; Yang et al., 2023). In addition to transferring knowledge between model architectures, this work also focuses on transferring knowledge between tasks.
@@
Task-oriented knowledge transfer from foundation models has been recently explored in the context of large langugae models (LLMs) by (Hsieh et al., 2023; Fu et al., 2023). These approaches use the rationales extracted from LLMs by chain-of-thought prompting to train small task-specific models. In this work, we focus on vision foundation models.
@@ A.2 Transfer sets
A.3 Task-specific heads
For classification tasks, the task-specific head is a linear classifier. The input to the classification layer is the final CLS token features for DINOv2 and OpenCLIP models, and the output of the final average pooling layer for MViT-V2 and FastViT models. For segmentation tasks, the task-specific head consists of a DeepLabV3 segmentation head (Chen et al., 2017) on top of the final spatial feature map and a spatial upsampling layer. As segmentation tasks requires high resolution features, we modify the image encoders such that their output spatial resolution is $1/8$ of the input image resolution. For DINOv2 and OpenCLIP models, we use a convolution stride of 8 instead of 14 for the patch embedding layer and resize the position embeddings accordingly. For MViT-V2 and FastViT models, we change the last two stride-2 convolutions to stride-1 convolutions.
A.4 Training details
@@ B.2 Crop-level retrieval from DataComp
For each crop in the gallery, we store two normalized features obtained by OpenCLIP’s (Ilharco et al., 2021) ViT-L/14 image encoders trained with OpenAI and Datacomp-1B pretraining datasets. For efficient retrieval, we use the cached feature store as in the dataset reinforcement strategy of Vasu et al. (2023b). For query images, we resize and center-crop to $224\times 224$ and run the same image encoders. For a pair of a gallery crop and a query image, we compute their similarity (for the $k$-NN search described in Section 3.4) by averaging the cosine similarity of their features computed by the two image encoders. We used this retrieval strategy to obtain retrieval-augmented transfer sets for the EuroSAT (Helber et al., 2019) query sets.
@@ B.6 Retrieval Visualizations
Figure 10 displays a t-SNE visualization of image features from the ADE20K dataset and a randomly sampled 10% of the YFCC15M dataset. Images from the ADE20K dataset occupy only a small region. This accounts for the performance drop in Figure 5(left) when retrieving more than 150K images.
@@ C.1 FastViT results
Tables 8, 7, 4, 5 and 6 present our experimental results on HAM10K, ADE20K, ImageNet, EuroSAT and Places365 datasets, respectively, in tabular format. The values in these tables correspond to the plots in Figures 4 and 11.
@@
Finetuned VFM (OpenCLIP) $74.72^{*}$ $83.31$ $84.98$ $86.05$
From scratch $14.46$ $56.35$ $68.90$ $75.17$
@@
DINO-Pretrain (CC3M) $33.29$ $63.89$ $71.36$ $75.97$
DINO-Pretrain (ImageNet) $52.27^{*}$ $70.64$ $74.79$ $77.41$
OpenCLIP Task-Agn $51.47^{*}$ $68.12$ $73.24$ $76.53$
(CC3M transfer) Task-Oriented $63.13$ $74.29$ $76.79$ $78.40$
@@
(ImageNet transfer) Task-Oriented $74.03$ $80.08$ $80.91$ $81.43$
Table 4: ImageNet classification accuracy for FastViT target model. The results marked with $*{}^{*}$ are obtained by training only the classification layer instead of the entire model in the finetuning stage. Full finetuning produced inferior results in these cases.
@@
Finetuned VFM (DINOv2) $47.56\pm 0.02$ $54.11\pm 0.27$ $56.95\pm 0.11$
Finetuned VFM (OpenCLIP) $48.30\pm 0.59$ $54.26\pm 0.19$ $57.03\pm 0.17$
IN-Pretrain $40.58\pm 0.14$ $47.96\pm 0.09$ $52.33\pm 0.14$
@@
DINO-Pretrain (CC3M) $40.29\pm 0.26$ $48.20\pm 0.16$ $52.74\pm 0.12$
DINO-Pretrain (Places365) $45.97\pm 0.18$ $50.85\pm 0.12$ $53.96\pm 0.06$
@@
(CC3M transfer) Task-Oriented $48.75\pm 0.01$ $53.33\pm 0.05$ $55.75\pm 0.06$
@@
(Places365 transfer) Task-Oriented $50.47\pm 0.01$ $54.82\pm 0.04$ $56.80\pm 0.02$
Table 6: Places365 classification accuracy for FastViT target model. The results marked with $*{}^{*}$ are obtained by training only the classification layer instead of the entire model in the finetuning stage. Full finetuning produced inferior results in these cases.

### v3

@@ 1 Introduction
Figure 1: Downstream task performance (EuroSAT classification) of FastViT target model with different pretraining approaches. Here, we finetune the pretrained models using 10 labeled training images per class. Task-oriented knowledge transfer from DINOv2 VFM with generic web data (CC3M) outperforms the popular ImageNet, CLIP and DINO pretraining approaches. Knowledge transfer with a transfer set curated using image retrieval performs significantly better than knowledge transfer with generic web data.
Currently, the computer vision community is witnessing the emergence of various vision and multi-modal foundation models pretrained on massive datasets (Oquab et al., 2023; Radford et al., 2021; Yuan et al., 2021; Li et al., 2023b; Wang et al., 2023b). These models have been shown to work well for many downstream tasks, especially, when task-specific labeled data is limited (Radford et al., 2021). However, they cannot be used for many resource-constrained applications due to their high inference computate cost. Also, many applications such as autonomous driving, medical image diagnostics, and industrial automation, focus on specific tasks or domains and need small task-specific models rather than a large Vision Foundation Model (VFM). This raises an important, timely and yet underexplored question: How can we leverage the knowledge from a large VFM to effectively train a small task-specific model for a new target task with limited labeled training data?
@@ 2 Approach and Contributions
We compare the proposed task-oriented knowledge transfer approach with task-agnostic knowledge transfer from VFMs, web-scale CLIP pretraining, supervised ImageNet pretraining, and self-supervised DINO (Caron et al., 2021) pretraining on five target tasks under limited labeled data settings  [^1] The target task data splits can be found at https://github.com/apple/ml-vfm-kt/tree/main. using two VFMs, namely DINOv2-ViT-L/14 (Oquab et al., 2023) and OpenCLIP-ViT-L/14 (Ilharco et al., 2021), and two target mobile architectures, namely FastViT-S12 (Vasu et al., 2023) and MobileViT-V2-1.0. We experiment with two transfer sets, a generic CC3M (Sharma et al., 2018) transfer set and a target task-related transfer set, and present several insightful findings to the community. To the best of our knowledge, there is no existing work that concretely establishes the below findings about leveraging VFMs for training small task-specific models with limited labeled data.
@@
  • Task-oriented knowledge transfer from VFMs outperforms the compute-intensive CLIP, widely-used supervised ImageNet and self-supervised DINO pretraining approaches by large margins (see Table 1), and is also computationally efficient (see Figures 1 and 4).
@@ 3.1 Experimental Setup
Alternative approaches: IN-Pretrain: Supervised pretraining on 1.28M labeled images from the ImageNet-1K dataset (Deng et al., 2009). CLIP-Pretrain: Contrastive language image pretraining on 0.7B image-text pairs from (Cao et al., 2023) following the affinity mimicking CLIP distillation approach of (Wu et al., 2023)222 (Wu et al., 2023) has shown that affinity mimicking distillation works better than the CLIP loss of (Radford et al., 2021). We use the best ViT-B/16 model from (Mehta et al., 2022) as teacher. DINO-Pretrain: Self-supervised pretraining on the unlabeled transfer set following the DINO approach of (Caron et al., 2021). We finetune the target model on labeled target task data after IN/CLIP/DINO pretraining.
@@
Target models: We use two recent efficient architectures, namely FastViT-S12 (Vasu et al., 2023) and MobileViT-V2-1.0 (Mehta & Rastegari, 2023), as image encoders for the target models. For brevity, we refer to them as FastViT and MViT-V2, respectively. We present the results for FastViT in this section and the results for MViT-V2 in Section C.2.
Task-specific heads: For classification tasks, the task-specific head is a linear classifier. The input to the classification layer is the final CLS token features for DINOv2 and OpenCLIP models, and the output of the final average pooling layer for MViT-V2 and FastViT models. For segmentation tasks, the task-specific head consists of a DeepLabV3 segmentation head (Chen et al., 2017) on top of the final spatial feature map and a spatial upsampling layer. As segmentation tasks requires high resolution features, we modify the image encoders such that their output spatial resolution is $1/8$ of the input image resolution. For DINOv2 and OpenCLIP models, we use a convolution stride of 8 instead of 14 for the patch embedding layer and resize the position embeddings accordingly. For MViT-V2 and FastViT models, we change the last two stride-2 convolutions to stride-1 convolutions.
@@ 3.2.1 Performance Improvement
When task-related transfer set is used, task-oriented transfer even outperforms finetuned VFM for some of the tasks when the amount of labeled data is small (see red and black curves in Figure 3)). This could be because the knowledge transfer process uses large-scale unlabeled target domain data that is not used while finetuning VFMs.
@@ 3.2.3 VFM Finetuning vs Linear Probing
Pretraining dataset - CC3M transfer set
ADE20K Number of labeled training images 1200 2401
Performance gain 2.19 2.03
ImageNet Percentage of labeled training images 10 25
Performance gain 3.8 5.8
HAM10K Maximum labeled training images per class 250 500
Performance gain 9.99 7.41
EuroSAT Number of labeled training images per class 5 10
Performance gain 3.3 4.9
Places365 Number of labeled training images per class 250 1000
Performance gain 0.47 0.76
Pretraining dataset - Target task-related transfer set
ADE20K Number of labeled training images 1200 2401
Performance gain 1.23 1.25
Table 3: Performance gain (for FastViT model) due to the final finetuning step of task-oriented knowledge transfer approach.
3.2.4 Contributions of Finetuning Step
The last step of the proposed task-oriented knowledge transfer approach is to finetune the target model on limited labeled data (see Figure 2(c)). One may wonder to what extent this finetuning step contributes to the performance, given that the target model learns to follow a VFM finetuned on the target task in the pretraining step. Table 3 shows the performance improvements due to the finetuning step. We see significant gains when generic CC3M transfer set is used during pretraining. When target task-related transfer set is used during pretraining, we see significant gains in the case of ADE20K segmentation task. We did not observe noticeable gains for the classification tasks in this case.
@@ 3.4 Retrieval Augmented Knowledge Transfer
Specifically, we use the limited target task labeled dataset as the query set $\mathcal{Q}$ and a large set of images sourced from web as the gallery $\mathcal{G}$. We use an encoder network $\phi$ to map all the images to a $d$-dimensional embedding space, and perform retrieval based on Euclidean distances in this space following a query balanced approach. For a query image $x_{q}\in\mathcal{Q}$, we define $k$-NN$(x_{q})$ to be the set of its $k$ nearest neighbors from $\mathcal{G}$. To retrieve $N$ images in total, we find the smallest $k$ for which $\bigcup_{x_{q}\in\mathcal{Q}}k$-NN$(x_{q})$ contains at least $N$ unique images. If $\bigcup_{x_{q}\in\mathcal{Q}}k$-NN$(x_{q})$ contains more than $N$ images, we drop the $k^{th}$ neighbor of randomly selected queries until the retrieved set has $N$ images. By giving equal weight to all the queries, this approach increases the diversity of the retrieved samples. We combine the retrieved set with the initial query set and use this retrieval augmented transfer set for task-oriented knowledge transfer.
@@
3.4.1 ADE20K Segmentation
Figure 5 (left) shows ADE20K segmentation performance of task-oriented knowledge transfer with retrieval augmented transfer sets of different sizes. Here, we use 4800 labeled images as both finetuning dataset and retrieval query set. The segmentation performance increases with the transfer set size until we reach 154K images and drops after that. Figure 6 displays a t-SNE visualization of image features from the ADE20K dataset and a randomly sampled 10% of the YFCC15M dataset. Images from the ADE20K dataset occupy only a small region suggesting that only a small subset of YFCC15M is relevant for ADE20K segmentation. Hence, as we increase the size of retrieved dataset, after certain point, low quality matches will be added to the transfer set resulting in performance drop.
@@ 3.4.1 ADE20K Segmentation
Using 154K as the target transfer set size, we curate different transfer sets by varying the number of query images used for retrieval. Table 4 (top) compares these curated transfer sets with the generic CC3M and the full ADE20K transfer sets. Retrieval augmented transfer sets clearly outperform both ADE20K and generic CC3M transfer sets.  [^4] The best performance in Figure 5(left) is lower than the performance for curated transfer set corresponding to 4800 query images in Table 4(top). This is because, we used shorter training runs (60K steps) to get the results in Figure 5(left), and once we identified the best transfer set size, we used longer training runs (200K steps) to get the results in Table 4(top).
@@ 4 Related Works
Knowledge distillation is a widely-used approach for transferring knowledge between model architectures by training one model to mimic the outputs of another model. Numerous distillation approaches have been proposed over the past decade based on various knowledge representations such as task logits (Hinton et al., 2015), intermediate features or embeddings (Heo et al., 2019; Tian et al., 2020), relations between samples (Park et al., 2019), attention maps (Zagoruyko & Komodakis, 2017), etc. Please refer to Wang & Yoon (2022); Hu et al. (2023) for an overview of existing distillation approaches. Some recent works have specifically focused on multi-modal distillation of image-language models (Fang et al., 2021; Li et al., 2023c; Wang et al., 2023a; Sun et al., 2023; Yang et al., 2023; Vasu et al., 2024). In addition to transferring knowledge between model architectures, this work also focuses on transferring knowledge between tasks.
@@
Task-agnostic knowledge transfer from VFMs has been explored in recent works such as DINOv2 (Oquab et al., 2023) and GSD (Huang et al., 2023). However, they did not evaluate the effectiveness of task-agnostic transfer in limited labeled data settings. Also, they use distillation approaches specifically designed for transformer architectures.
Task-oriented knowledge transfer has been recently explored in the context of large langugae models (LLMs) by (Hsieh et al., 2023; Fu et al., 2023). These approaches use the rationales extracted from LLMs by chain-of-thought prompting to train small task-specific models. In this work, we focus on vision foundation models.
@@ 5 Conclusions
Future work: In this work, we only used labeled target task data to finetune VFMs. A potential future work is to leverage additional large-scale unlabeled data to better adapt VFMs to the target task/domain, thereby eventually improving the small task-specific model trained with knowledge transfer from VFMs. However, finetuning VFMs on large scale datasets could be computationally expensive. Another potential future work is to explore more sophisticated transfer set curation strategies based on active learning and adversarial filtering that may result in better transfer sets that lead to better target model performance.
@@
Acknowledgements
We thank Ting-Yao Hu, Karren Yang, Mohammad Samragh Razlighi, Pavan Kumar Anasosalu Vasu, and Barry Theobald from Apple for their valuable suggestions.
@@ A.1 Target Task Datasets
The training, validation and test splits used in this work can be found at https://github.com/apple/ml-vfm-kt/tree/main.
@@ A.2 Transfer Sets
A.3 Training Details
@@ B.2 Crop-level Retrieval from DataComp
For each crop in the gallery, we store two normalized features obtained by OpenCLIP’s (Ilharco et al., 2021) ViT-L/14 image encoders trained with OpenAI and Datacomp-1B pretraining datasets. For efficient retrieval, we use the cached feature store as in the dataset reinforcement strategy of Vasu et al. (2024). For query images, we resize and center-crop to $224\times 224$ and run the same image encoders. For a pair of a gallery crop and a query image, we compute their similarity (for the $k$-NN search described in Section 3.4) by averaging the cosine similarity of their features computed by the two image encoders. We used this retrieval strategy to obtain retrieval-augmented transfer sets for the EuroSAT (Helber et al., 2019) query sets.
@@ C.1 FastViT Results
Tables 9, 8, 5, 6 and 7 present our experimental results on ImageNet, EuroSAT, Places365, ADE20K and HAM10K datasets, respectively, in tabular format. The values in these tables correspond to the plots in Figures 3 and 11.
@@
Finetuned VFM (OpenCLIP) $74.72^{*}$ $83.31^{\hskip 3.0pt}$ $84.98$ $86.05$
From scratch $14.46^{\hskip 3.0pt}$ $56.35^{\hskip 3.0pt}$ $68.90$ $75.17$
@@
DINO-Pretrain (CC3M) $33.29^{\hskip 3.0pt}$ $63.89^{\hskip 3.0pt}$ $71.36$ $75.97$
DINO-Pretrain (ImageNet) $52.27^{*}$ $70.64^{\hskip 3.0pt}$ $74.79$ $77.41$
OpenCLIP Task-Agn $51.47^{*}$ $68.12^{\hskip 3.0pt}$ $73.24$ $76.53$
(CC3M transfer) Task-Oriented $63.13^{\hskip 3.0pt}$ $74.29^{\hskip 3.0pt}$ $76.79$ $78.40$
@@
(ImageNet transfer) Task-Oriented $74.03^{\hskip 3.0pt}$ $80.08^{\hskip 3.0pt}$ $80.91$ $81.43$
Table 5: ImageNet classification accuracy for FastViT target model. The results marked with are obtained by training only the classification layer instead of the entire model in the finetuning stage. Full finetuning produced inferior results in these cases.
@@
Finetuned VFM (DINOv2) $47.56\pm 0.02^{\hskip 3.0pt}$ $54.11\pm 0.27$ $56.95\pm 0.11$
Finetuned VFM (OpenCLIP) $48.30\pm 0.59^{\hskip 3.0pt}$ $54.26\pm 0.19$ $57.03\pm 0.17$
IN-Pretrain $40.58\pm 0.14^{\hskip 3.0pt}$ $47.96\pm 0.09$ $52.33\pm 0.14$
@@
DINO-Pretrain (CC3M) $40.29\pm 0.26^{\hskip 3.0pt}$ $48.20\pm 0.16$ $52.74\pm 0.12$
DINO-Pretrain (Places365) $45.97\pm 0.18^{\hskip 3.0pt}$ $50.85\pm 0.12$ $53.96\pm 0.06$
@@
(CC3M transfer) Task-Oriented $48.75\pm 0.01^{\hskip 3.0pt}$ $53.33\pm 0.05$ $55.75\pm 0.06$
@@
(Places365 transfer) Task-Oriented $50.47\pm 0.01^{\hskip 3.0pt}$ $54.82\pm 0.04$ $56.80\pm 0.02$
Table 7: Places365 classification accuracy for FastViT target model. The results marked with are obtained by training only the classification layer instead of the entire model in the finetuning stage. Full finetuning produced inferior results in these cases.